# Rocky planet formation in compact disks around M dwarfs

### v2

@@ 4.3 Changing the gas disk viscosity and pebble flux: PA-M01-high-icy
In this scenario we considered a central star of M M⊙, embryos located outside the snowline growing by pebble accretion in a gas disk with a higher viscosity:  and also a lower pebble flux F. In this case the sample of embryos migrated inwards much slower than in the scenarios with a lower viscosity. The resulting planets located close to the star ( au) accreted pebbles while migrating inwards during the whole disk lifetime. In this scenario, in all simulations there is a sample of small planets located at further distances from the star up to almost 2 au (initial location of the outermost planet). While the close-in planets remind in quasi-circular orbits, the outermost planets present eccentricities . The resulting planets present fewer collisions, as just less than 30 of the embryos collided among them and no one collided with the central star, in any of the simulations. In all the simulations, the collision among embryos occurred both before and after the dissipation of the gas disk at 10 Myr. The latest collisions occurred at around 45 Myr. Thus, the resulting planets achieved the final masses and configurations at around 45 Myr.
@@ 4.4 Changing the accretion mechanism: PLA-M01-low-icy
In this scenario we considered a star of M M⊙, a low gas disk viscosity () and a sample of embryos located beyond the snowline that grow by planetesimal accretion, which were distributed both inside and outside the snowline. The resulting planets were also located close-in to the star ( au) but in this case there was also a tail of planetesimals at further distances (). The planets present compact orbital configurations in quasi-circular, a co-planar orbits. In this case, the embryos also suffered a fast inward migration, where most of the collision among them and with planetesimals occurred within 300,000 yr.
@@
The final masses of the resulting planets are between 0.15 and 3 M⊕. In this case, the final masses are smaller than when considering embryos growing by pebble accretion in a gas disk with low viscosity. In this scenario, more planets can be formed per systems, as in most of the simulations each system has either 5 or 6 planets. Roughly, half of the planets in each system present masses less than 1 M⊕. Such masses always present between 45 and 50 of water, as even though the planets accreted a bigger number of dry planetesimals, the major contribution was of water-rich embryos together with some water-rich planetesimals.
@@ 4.5 Changing the stellar mass: M (M03-low-icy) and M (M05-low-icy)
Pebble accretion
As in our standard model, we consider a sample of embryos located outside the snowline growing by pebble accretion in a gas disk with low viscosity. When we consider a star of M M⊙, the sample of embryos migrated inwards within just 70,000 yr, due to mostly the fact that they are immersed in a more massive gas disk (gas density a factor 2 higher than in our standard model). During this time, the outermost planet reached its isolation mass. The planetary systems have between 2 and 5 planets, each of them in compact configurations ( au), in quasi-circular and co-planar orbits (except just one simulation in which no planet survived). In all the simulations, 90 of the embryos collided among them, mostly within 70,000 yr. However, in 60 of the simulations, few outer embryos collided among them just after the gas dissipated from the disk at around 10 Myr. Thus, after that time, they reached their final configurations and masses. Regarding the collisions with the star, just in 1 simulation, 1 embryo collided with the star.
@@
Planetesimal accretion
In the scenarios in which we considered that the sample of embryos was growing by planetesimal accretion instead of growing by pebble accretion, the simulated planets present almost the same range in semi-major axis and number of planets per system. They are also in compact configurations and in quasi-circular and co-planar orbits. The main difference is that there are smaller planets in each system, which are mostly the outermost planets. Thus, the planetesimal accretion allows the formation of smaller planets, that can not be formed if we consider pebble accretion.
For a star of 0.3 M⊙ as for the star of 0.5 M⊙, the embryos suffer a slower inward migration as the majority of the collision between embryos and planetesimal occurred within the first 200,000 yr. In most of the simulations some embryos collided among them during the gas disk but just in 30 of the simulations there were late collisions between embryos after the gas disk dissipated and up to 40 Myr. The final masses of the resulting planets are analogous as when considering pebble accretion, but in this case planets smaller than 2 M⊕ could be formed as well. As for a star of 0.1 M⊙, the planets present almost 50 of their mass in water. Whereas the smallest planets truly exist or not can not be confirmed yet with the current instrumental sensitivity.
@@ 4.6 Comparison among the formation scenarios
We conclude that for a star of 0.1 M⊙, we can form multiple close-in Earth-like and some super-Earth like planets both by pebble or planetesimal accretion in a gas disk with low viscosity (). The main difference is that when considering planetesimal accretion all the planets are water worlds, and we can form also smaller planets. Additionally, changing the initial embryo distribution either inside or outside the snowline has an impact just in the final amount of water in mass. We can form dry worlds in close-in orbits, just if we consider embryos initially located inside the snowline. Moreover, if we consider a gas disk with a high viscosity () and lower pebble flux just planets with masses less than 0.5 M⊕ can be formed, as the migration rate is slower than in a disk with lower viscosity, and the available pebble mass seems not enough to form Earth-like planets.
@@
We highlight that in the scenarios in which we let a sample of embryos growth by pebble accretion in a disk with low viscosity (), the pebble accretion was efficient. For a star of 0.1 M⊙ the average of pebble efficiency was around 30, while for a star of 0.3 M⊙ and 0.5 M⊙, in each case, the average in efficiency was around 60. In Figure 6 we show the semi-major axis and the initial planetary masses of the embryos, the planetary masses at 10 Myr (gas disk lifetime) and the final planetary masses at 50 Myr (integration time) of the simulated planets of the formation scenarios that included pebble accretion in a gas disk with low viscosity. We compare the masses and location of the planets with the isolation masses and the location of the inner edge of the disk for each stellar mass. Despite the fact that the isolation mass can reach 6 and 10 M⊕ in the outer disk, depending on the stellar mass, the maximum isolation mass that each individual embryo could reach was between 1 and 3 M⊕, as they rapidly migrated inwards. We can see how the final planetary mass of the outer planets is the same as the isolation mass, but the inner planets have final masses between 2 to 5 times bigger than the isolation mass in the inner regions. This is mostly due to collisions among embryos in the inner disk during the gas disk lifetime, and in some cases collisions after the gas dissipated from the disk. We conclude that the final planetary masses are mostly determined by the amount of collisions between embryos due to the gravitational interactions between the bodies. Moreover, we found that the mass of super-Earth planets increase almost linearly with the stellar mass, in agreement with previous works (e.g. Liu et al., 2019).
@@ 5 Simulated planetary systems vs observed low-mass exoplanets around M dwarfs
In this section we compare the resulting simulated planetary systems from each formation scenario with the sample of confirmed low-mass close-in exoplanets (M M⊕ and au) around M dwarfs 111https://exoplanetarchive.ipac.caltech.edu/, in terms of masses, semi-major axis and orbital periods. We just considered the sample of planets with values of planetary masses, which are 112 exoplanets up to date. From the total amount of exoplanets, 62 were discovered with radial velocity and 38 with transit technique, with planetary masses estimated by transit timing variations in this last case. They were principally discovered with CARMENES (Quirrenbach et al., 2014), HARPS (Mayor et al., 2011) and TESS (Ricker et al., 2015), among other instruments.
@@ 5.1 Planetary systems with M M⊙
We focus first on the final simulated planets of the scenarios in which we assumed a star of M M⊙, with the sample of low-mass close-in exoplanets around stars  that have an estimation of their masses. In Figure 7 we show the planetary masses and semi-major axis of the simulated planets and the confirmed exoplanets. We also show the isolation mass for a star of 0.1 M⊙ at 1 Myr (as the planets migrated inwards early in their evolution) and the detection limit in mass for planets around M dwarfs (0.2 M⊕). Additionally, in Figure 8 we compare the histograms in planetary mass and semi-major axis of the resulting planets of each scenario (around 40 planets each) with the confirmed exoplanet sample (35 planets). In all the scenarios, we generate planets with semi-major axis close-in to the star, which match the observations. On the other hand, as it was explained in Section 4, the scenarios in which we considered a sample of embryos that grow by pebble accretion in a gas disk with low viscosity () present the planets with higher masses in compact configurations. Thus, there is an overlap between the simulated and observed planets for masses less than 5 M⊕ as it can be seen in Figure 8. Finally, the scenario in which the embryos grow by planetesimal accretion formed a bigger amount of small planets ( M⊕), while the scenario with a gas disk with a higher viscosity () present just small planets ( M⊕), many of them below the detection limit (see Figure 6).
@@ 5.2 Planetary systems formed by pebble accretion: M M⊙, M M⊙, and M M⊙
We compared the simulated planets formed in the scenarios that included pebble accretion in a gas disk with low viscosity () with the exoplanets around M dwarfs, as such scenarios are the most efficient in forming planets bigger than 1 M⊕ (see Section 4). In Figure 9 we show the planetary mass and semi-major axis of the simulated planets and the confirmed low-mass close-in exoplanets around M dwarfs of different masses. We can see how the simulated planets and the observed exoplanets present, in most of the cases, masses higher than the isolation mass. As we discuss in the previous section, the final masses could be achieved mostly from planet-planet collisions, specially for higher stellar masses. In addition, in Figure 10 we show the histograms in planetary mass and semi-major axis of the simulated planets in each scenario (around 30 planets each) with three different subsamples of exoplanets, each of them associated to different stellar masses:  (35 planets),  (39 planets) and  (38 planets). In each scenario, we generate planets close-in to the star that match the observations. In particular, regarding the planetary masses, the simulated planets from the scenario with a star of 0.1 M⊙ overlap the exoplanets with masses between 0.7 and 5 M⊕, the ones from the scenario with a mass of 0.3 M⊙ overlap the exoplanets with masses between 2 and 10 M⊕ and the ones from the scenario with 0.5 M⊙ overlap the exoplanets with masses between 3 and 17 M⊕.
@@
Secondly, we compared the simulated planetary systems with the observed systems with more than one planet detected. In Figure 11 we show the planetary architectures (mass and semi-major axis) of the simulated planets of the scenarios in which we include pebble accretion and low gas disk viscosity, together with each observed planetary system around stars of different masses: ,  and . We can see how the simulated planetary systems and the observed planetary systems share the same range of semi-major axis, planetary masses and number of planets per mass (see Table 3). Also, we can see the agreement in the fact that the innermost planets are located further from the star when increasing the stellar mass. In particular, for a star of 0.1 M⊙ we could reproduce planetary systems analogous to Trappist-1 in 20 of the simulations. For a star of 0.3 M⊙ we can reproduce some of the systems with 2 or 3 planets with analogous masses to the observed exoplanets, but in more compact configurations. Lastly, for a star of 0.5 M⊙ we can reproduce few of the systems in terms of masses and multiplicity, but also in more compact configurations. We highlight that in this case we can see simulated planets with masses higher than 10 M⊕ that so far are not been detected as part of multi-planetary systems, but as single planets. This could be because of biases in radial velocity and transit techniques in the observed planetary systems. We could be missing smaller planets at further distances from the star (transit technique bias) or smallest planets in between more massive planets (radial velocity technique bias) that can not be detected.
@@
On the other hand, most of the simulated planets with masses 5 M⊕ present less amount of water in mass than the observed ones. In this case, to have a good contrast with observations, it will be crucial to take into account the influence of planetary atmosphere in the water content (see Section 6 for further discussion).
@@
Additionally, we compared the period ratio of adjacent simulated planets and confirmed exoplanets in multiple systems. In Figure 13 we show the cumulative distributions of the period ratio of adjacent planets in each scenario that include pebble accretion and a low gas disk viscosity with the observed exoplanet sample around stars with similar masses. As it can be seen also from Figure 11, many of the simulated planetary systems are in more compact configurations than the observed planetary systems, as the period ratio of the simulated planetary systems are on average 2 times lower than the observed ones. The simulated systems present period ratios less than 3 while the observed systems present period ratios up to 8. We suggest that this discrepancy may be due to the fact that there are missing planets in between the confirmed exoplanets, at least for the case of a star of 0.1 M⊙. For example, up to date is not possible to detect planets with masses less than 3 M⊕ with orbital periods between 10 and 40 days (or with semi-major axis between 0.04 and 0.1 au) (Ribas et al., 2023). For the more massive M dwarfs, which host more massive planets, it could be that our model is predicting more compact orbits because we are not considering later collisions among the planets at a Gyr timescale. This is something to be explored in future works.
@@
From Figure 13 we could also infer that in the simulated systems there are some pairs of planets that are in commensurable orbits: . Each pair of such adjacent planets satisfied that , with  and  the orbital period of each adjacent planet and  and  two integer numbers. In the scenario with a star of 0.1 M⊙, 20 of the planets are in commensurability 3:2 and 10 in 2:1. In the scenario with a star of 0.3 M⊙, 15 of the planets are in 3:2 and almost 10 in 5:3. Lastly, in the scenario with a star of 0.5 M⊙ just less than 10 of the planets are in 2:1. In most of the cases, such commensurability was reached after the last collisions among embryos. It happened earlier in the gas phase for a star of 0.1 M⊙ and once the gas dissipated from the disk for the more massive M dwarfs. We can see also how there are more planets in commensurable orbits around the less massive M dwarfs, and the number is decreasing towards the more massive ones. This tendency is also seen in the observed exoplanet samples, even though it seems that our simulations are overestimating the number of planets in commensurable orbits. This trend seems to be in agreement with the lack of resonances found around Sun-like stars regarding the Kepler sample, which is also shown from a theoretical perspective when modelling super-Earths around Sun-like stars (e.g. Izidoro et al., 2017, 2021) .
@@
Finally, in Figure 13 we also show the cumulative distributions of the period of the innermost planets in each scenario that include pebble accretion and a low gas disk viscosity and the related observed exoplanet sample. For each stellar mass, we obtained simulated innermost planets with periods as low as the confirmed exoplanet sample. For a star of 0.1 M⊙ we could form innermost planets with periods less than 3 days, however, we are not able to reproduce the ones with periods between 3 and 10 days (or semi-major axis between 0.02 and 0.04 au). For a star of 0.3 M⊙ we could form innermost planets with periods 3 to 12 days, which is in agreement with observations, but we are underestimating the planets with period less than 5 days (or semi-major axis less than 0.04 au). Lastly, for a star of 0.5 M⊙ we could form innermost planets with periods up to 12 days, but we are not able to reproduce the innermost planets with periods from 12 to 20 days (or semi-major axis between 0.07 and 0.11 au). We are also overestimating the planets with periods less than 12 days. Even though our simulated sample is not matching perfectly with the observations, the difference in between the innermost simulated and observed planets’ periods is the only few days or between 0.02 and 0.04 au in semi-major axis. We think that such differences are due to the selection of the initial stellar rotation period that sets the inner edge of the disk. We claim that it could be possible to better reproduce the observations if we assumed different values of the stellar rotation period, given the range in the estimations of the rotational period of the stars, but this is out of the scope of this work.

### v3

@@ 4.3 Changing the gas disk viscosity and pebble flux: PA-M01-high-icy
In this scenario we considered a central star of M M⊙, embryos located outside the snowline growing by pebble accretion in a gas disk with a higher viscosity: and also a lower pebble flux F. In this case the sample of embryos migrated inward much slower than in the scenarios with a lower viscosity. The resulting planets located close to the star ( au) accreted pebbles while migrating inward during the whole disk lifetime. In this scenario, in all simulations there is a sample of small planets located at further distances from the star up to almost 2 au (initial location of the outermost planet). While the close-in planets remind in quasi-circular orbits, the outermost planets present eccentricities . The resulting planets present fewer collisions, as just less than 30 of the embryos collided among them and no one collided with the central star, in any of the simulations. In all the simulations, the collision among embryos occurred both before and after the dissipation of the gas disk at 10 Myr. The latest collisions occurred at around 45 Myr. Thus, the resulting planets achieved the final masses and configurations at around 45 Myr.
@@ 4.4 Changing the accretion mechanism: PLA-M01-low-icy
In this scenario we considered a star of M M⊙, a low gas disk viscosity (), and a sample of embryos located beyond the snowline that grow by planetesimal accretion, which were distributed both inside and outside the snowline. The resulting planets were also located close-in to the star ( au) but in this case there was also a tail of planetesimals at further distances (). The planets present compact orbital configurations in quasi-circular, a co-planar orbits. In this case, the embryos also suffered a fast inward migration, where most of the collision among them and with planetesimals occurred within 300,000 yr.
@@
The final masses of the resulting planets are between 0.15 and 3 M⊕. In this case, the final masses are smaller than when considering embryos growing by pebble accretion in a gas disk with low viscosity. In this scenario, more planets can be formed per systems, as in most of the simulations each system has either 5 or 6 planets. Roughly, half of the planets in each system present masses of less than 1 M⊕. Such masses are always between 45 and 50 water; even though the planets accreted a greater number of dry planetesimals, the major contribution was from water-rich embryos together with some water-rich planetesimals.
@@ 4.5 Changing the stellar mass: M (M03-low-icy) and M (M05-low-icy)
4.5.1 Pebble accretion
As in our standard model, we consider a sample of embryos located outside the snowline growing by pebble accretion in a gas disk with low viscosity. When we consider a star of M M⊙, the sample of embryos migrated inward within just 70,000 yr, due to mostly the fact that they are immersed in a more massive gas disk (gas density a factor 2 higher than in our standard model). During this time, the outermost planet reached its isolation mass. The planetary systems have between 2 and 5 planets, each of them in compact configurations ( au), in quasi-circular and co-planar orbits (except just one simulation in which no planet survived). In all the simulations, 90 of the embryos collided among them, mostly within 70,000 yr. However, in 60 of the simulations, few outer embryos collided among them just after the gas dissipated from the disk at around 10 Myr. Thus, after that time, they reached their final configurations and masses. Regarding the collisions with the star, just in 1 simulation, 1 embryo collided with the star.
@@ 4.5.1 Pebble accretion
4.5.2 Planetesimal accretion
In the scenarios in which we let the sample of embryos grow by planetesimal accretion instead of by pebble accretion, the simulated planets present almost the same range in semi-major axis and number of planets per system. They are also in compact configurations and in quasi-circular and co-planar orbits. The main difference is that there are smaller planets in each system, which are mostly the outermost planets. Thus, the planetesimal accretion allows the formation of smaller planets, which cannot be formed if we consider pebble accretion.
For a star of 0.3 M⊙ as for the star of 0.5 M⊙, the embryos suffered a slower inward migration as the majority of the collision between embryos and planetesimal occurred within the first 200,000 yr. In most of the simulations some embryos collided among them during the gas disk, but just in 30 of the simulations there were late collisions between embryos after the gas disk dissipated and up to 40 Myr. The final masses of the resulting planets are analogous as when considering pebble accretion, but in this case planets smaller than 2 M⊕ could be formed as well. As for a star of 0.1 M⊙, the planets present almost 50 of their mass in water. Whereas the smallest planets truly exist cannot be confirmed yet with the current instrumental sensitivity.
@@ 4.6 Comparison among the formation scenarios
We conclude that for a star of 0.1 M⊙, we can form multiple close-in Earth-like and some super-Earth-like planets both by pebble or planetesimal accretion in a gas disk with low viscosity (). The main difference is that when considering planetesimal accretion, all the planets are water worlds, and we can form also smaller planets. Additionally, changing the initial embryo distribution either inside or outside the snowline has an impact just in the final amount of water in mass. We can form dry worlds in close-in orbits, just if we consider embryos initially located inside the snowline. Moreover, if we consider a gas disk with a high viscosity () and lower pebble flux just planets with masses less than 0.5 M⊕ can be formed, as the migration rate is slower than in a disk with lower viscosity, and the available pebble mass seems not enough to form Earth-like planets.
@@
We highlight that in the scenarios in which we let a sample of embryos grow by pebble accretion in a disk with low viscosity (), the pebble accretion was efficient. For a star of 0.1 M⊙ the average of pebble efficiency was around 30, while for a star of 0.3 M⊙ and 0.5 M⊙, in each case, the average in efficiency was around 60. In Figure 6 we show the semi-major axis and the initial planetary masses of the embryos, the planetary masses at 10 Myr (gas disk lifetime) and the final planetary masses at 50 Myr (integration time) of the simulated planets of the formation scenarios that included pebble accretion in a gas disk with low viscosity. We compare the masses and location of the planets with the isolation masses and the location of the inner edge of the disk for each stellar mass. Despite the fact that the isolation mass can reach 6 and 10 M⊕ in the outer disk, depending on the stellar mass, the maximum isolation mass that each individual embryo could reach was between 1 and 3 M⊕, as they rapidly migrated inward. We can see how the final planetary mass of the outer planets is the same as the isolation mass, but the inner planets have final masses between two and five times bigger than the isolation mass in the inner regions. This is mostly due to collisions among embryos in the inner disk during the gas disk lifetime, and in some cases collisions after the gas dissipated from the disk. We conclude that the final planetary masses are mostly determined by the amount of collisions between embryos due to the gravitational interactions between the bodies. Moreover, we found that the mass of super-Earth planets increase almost linearly with the stellar mass, in agreement with previous works (e.g., Liu et al., 2019).
@@ 5 Simulated planetary systems vs observed low-mass exoplanets around M dwarfs
In this section we compare the resulting simulated planetary systems from each formation scenario with the sample of confirmed low-mass close-in exoplanets (M M⊕ and au) around M dwarfs,111https://exoplanetarchive.ipac.caltech.edu/ in terms of masses, semi-major axis and orbital periods. We just considered the sample of planets with values of planetary masses, which are 112 exoplanets to date. From the total amount of exoplanets, 62 were discovered with radial velocity and 38 with transit technique, with planetary masses estimated by transit timing variations in this last case. They were principally discovered with CARMENES (Quirrenbach et al., 2014), HARPS (Mayor et al., 2011) and TESS (Ricker et al., 2015), among other instruments.
@@ 5.1 Planetary systems with M M⊙
We focus first on the final simulated planets of the scenarios in which we assumed a star of M M⊙, with the sample of low-mass close-in exoplanets around stars that have an estimation of their masses. In Figure 7 we show the planetary masses and semi-major axis of the simulated planets and the confirmed exoplanets. We also show the isolation mass for a star of 0.1 M⊙ at 1 Myr (as the planets migrated inward early in their evolution) and the detection limit in mass for planets around M dwarfs (0.2 M⊕). Additionally, in Figure 8 we compare the histograms in planetary mass and semi-major axis of the resulting planets of each scenario (around 40 planets each) with the confirmed exoplanet sample (35 planets). In all the scenarios, we generated planets with semi-major axis close-in to the star, which match the observations. On the other hand, as it was explained in Section 4, the scenarios in which we considered a sample of embryos that grow by pebble accretion in a gas disk with low viscosity () present the planets with higher masses in compact configurations. Thus, there is an overlap between the simulated and observed planets for masses less than 5 M⊕ as it can be seen in Figure 8. Finally, the scenario in which the embryos grow by planetesimal accretion formed a bigger amount of small planets ( M⊕), while the scenario with a gas disk with a higher viscosity () present just small planets ( M⊕), many of them below the detection limit (see Figure 6).
@@ 5.2 Planetary systems formed by pebble accretion: M M⊙, M M⊙, and M M⊙
We compared the simulated planets formed in the scenarios that included pebble accretion in a gas disk with low viscosity () with the exoplanets around M dwarfs, as such scenarios are the most efficient in forming planets bigger than 1 M⊕ (see Section 4). In Figure 9 we show the planetary mass and semi-major axis of the simulated planets and the confirmed low-mass close-in exoplanets around M dwarfs of different masses. We can see how the simulated planets and the observed exoplanets present, in most of the cases, masses higher than the isolation mass. As we discuss in the previous section, the final masses could be achieved mostly from planet-planet collisions, especially for higher stellar masses. In addition, in Figure 10 we show the histograms in planetary mass and semi-major axis of the simulated planets in each scenario (around 30 planets each) with three different subsamples of exoplanets, each of them associated with different stellar masses: (35 planets), (39 planets), and (38 planets). In each scenario, we generate planets close-in to the star that match the observations. In particular, regarding the planetary masses, the simulated planets from the scenario with a star of 0.1 M⊙ overlap the exoplanets with masses between 0.7 and 5 M⊕, the ones from the scenario with a mass of 0.3 M⊙ overlap the exoplanets with masses between 2 and 10 M⊕, and the ones from the scenario with 0.5 M⊙ overlap the exoplanets with masses between 3 and 17 M⊕.
@@
Secondly, we compared the simulated planetary systems with the observed systems with more than one planet detected. In Figure 11 we show the planetary architectures (mass and semi-major axis) of the simulated planets of the scenarios in which we include pebble accretion and low gas disk viscosity, together with each observed planetary system around stars of different masses: , and . We can see how the simulated planetary systems and the observed planetary systems share the same range of semi-major axis, planetary masses and number of planets per mass (see Table 3). We can also see the agreement in the fact that the innermost planets are located further from the star when increasing the stellar mass. In particular, for a star of 0.1 M⊙ we could reproduce planetary systems analogous to Trappist-1 in 20 of the simulations. For a star of 0.3 M⊙ we can reproduce some of the systems with two or three planets with analogous masses to the observed exoplanets, but in more compact configurations. Lastly, for a star of 0.5 M⊙ we can reproduce a few of the systems in terms of mass and multiplicity, but also in more compact configurations. We highlight that in this case we can see simulated planets with masses higher than 10 M⊕ that so far are not been detected as part of multi-planetary systems, but as single planets. This could be because of biases in radial velocity and transit techniques in the observed planetary systems. We could be missing smaller planets at further distances from the star (transit technique bias) or small planets between more massive planets (radial velocity technique bias) that cannot be detected.
@@
On the other hand, most of the simulated planets with masses 5 M⊕ present less water in mass than the observed planets. In this case, to have a good contrast with observations, it will be crucial to take into account the influence of planetary atmosphere in the water content (see Section 6 for further discussion).
@@
Additionally, we compared the period ratio of adjacent simulated planets and confirmed exoplanets in multiple systems. In Figure 13 we show the cumulative distributions of the period ratio of adjacent planets in each scenario that include pebble accretion and a low gas disk viscosity with the observed exoplanet sample around stars with similar masses. As it can be seen also from Figure 11, many of the simulated planetary systems are in more compact configurations than the observed planetary systems, as the period ratio of the simulated planetary systems are on average two times lower than the observed ones. The simulated systems present period ratios of less than 3, while the observed systems present period ratios up to 8. We suggest that this discrepancy may be due to the fact that there are missing planets between the confirmed exoplanets, at least for the case of a star of 0.1 M⊙. For example, currently it is not possible to detect planets with masses less than 3 M⊕ with orbital periods between 10 and 40 days (or with semi-major axis between 0.04 and 0.1 au) (Ribas et al., 2023). For the more massive M dwarfs, which host more massive planets, it could be that our model is predicting more compact orbits because we are not considering later collisions among the planets at a Gyr timescale. This is something to be explored in future works.
@@
From Figure 13 we could also infer that in the simulated systems there are some pairs of planets that are in commensurable orbits: . Each pair of such adjacent planets satisfied that , with and the orbital period of each adjacent planet and and two integer numbers. In the scenario with a star of 0.1 M⊙, 20 of the planets are in commensurability 3:2 and 10 in 2:1. In the scenario with a star of 0.3 M⊙, 15 of the planets are in 3:2 and almost 10 in 5:3. Lastly, in the scenario with a star of 0.5 M⊙ just less than 10 of the planets are in 2:1. In most of the cases, such commensurability was reached after the last collisions among embryos. It happened earlier in the gas phase for a star of 0.1 M⊙ and once the gas dissipated from the disk for the more massive M dwarfs. We can see also how there are more planets in commensurable orbits around the less massive M dwarfs, and the number is decreasing toward the more massive ones. This tendency is also seen in the observed exoplanet samples, even though it seems that our simulations are overestimating the number of planets in commensurable orbits. This trend seems to be in agreement with the lack of resonances found around Sun-like stars regarding the Kepler sample, which is also shown from a theoretical perspective when modeling super-Earths around Sun-like stars (e.g., Izidoro et al., 2017, 2021) .
@@
Finally, in Figure 13 we also show the cumulative distributions of the period of the innermost planets in each scenario that include pebble accretion and a low gas disk viscosity and the related observed exoplanet sample. For each stellar mass, we obtained simulated innermost planets with periods as low as the confirmed exoplanet sample. For a star of 0.1 M⊙ we could form innermost planets with periods less than 3 days, however, we are not able to reproduce the ones with periods between 3 and 10 days (or semi-major axis between 0.02 and 0.04 au). For a star of 0.3 M⊙ we could form innermost planets with periods 3 to 12 days, which is in agreement with observations, but we are underestimating the planets with period less than 5 days (or semi-major axis less than 0.04 au). Lastly, for a star of 0.5 M⊙ we could form innermost planets with periods up to 12 days, but we are not able to reproduce the innermost planets with periods from 12 to 20 days (or semi-major axis between 0.07 and 0.11 au). We are also overestimating the planets with periods less than 12 days. Even though our simulated sample does not match perfectly with the observations, the difference between the innermost simulated and observed planets’ periods is the only few days or between 0.02 and 0.04 au in semi-major axis. We think that such differences are due to the selection of the initial stellar rotation period that sets the inner edge of the disk. We claim that it could be possible to better reproduce the observations if we assumed different values of the stellar rotation period, given the range in the estimations of the rotational period of the stars, but this is out of the scope of this work.